In a year a man smokes about 2.5kg of tobacco at 75p for 50g and 6000 cigarettes costing 31p for 20. What is his average weekly expenditure on smoking?
Answer by Boreal(15235)

In a year, the cost of tobacco is 50/75=2500/x
50x=2500*75
x=50*75=3750p
==============
6000/20=x/31
x=300*31=9300p
Total is 13,050p
52 weeks in a year
divide 13050/52=250.96 or 251 p per week.
=================
Check
in a week, just over 48 gm of tobacco or nearly 75 p
in a week, just over 115 cigarettes or nearly 186 p
that adds to 261 p, above the answer consistent with the two parts being slight overestimates.